OpenStudy (anonymous):

Find the midpoint of the segment with the given endpoints. M(–2, 1) and O(–3, 2)

OpenStudy (anonymous):

is it -0.5, -0.5?

OpenStudy (anonymous):

i thin the second ones a polstive

OpenStudy (anonymous):

A. (–2.5, 1.5) B. (–0.5, –0.5) C. (0.5, –0.5) D. (–3, –5)

OpenStudy (anonymous):

i can NOT get this answer wrong

OpenStudy (anonymous):

Well how are you going about getting the midpoint? Something like this? (X2-X1)/2,(Y2-Y1)/2

OpenStudy (anonymous):

A

OpenStudy (amistre64):

a midpoint is the average of the other 2 points add em up and divide by 2

OpenStudy (anonymous):

Rather (X2+X1)/2,(Y2+Y1)/2...which is indeed the average.

OpenStudy (anonymous):

answers A pertty shure

OpenStudy (anonymous):

so i got -2, 1.5

OpenStudy (amistre64):

-2 aint halfway between -2 and -3 :) but you mighta meant -2.5

OpenStudy (anonymous):

(-2+-3)/2 = -5/2 = -2.5 (2+1)/2 = 3/2 = 1.5 Make sense?

OpenStudy (anonymous):

sorry i meant -2.5 lol